OpenStudy (anonymous):

5y + 2 = 1/2 (10y + 4)

OpenStudy (texaschic101):

5y + 2 = 1/2(10y + 4) --- distribute through the parenthesis 5y + 2 = 5y + 2 5y - 5y = 2 - 2 0 = 0 This equation has INFINITE SOLUTIONS